OpenStudy (anonymous):

whats x in 2=ln(x+9)

OpenStudy (anonymous):

symmetrical across x=0 (this is true for all functions that contain only even powers of x). domain: x^2-9>0 x^2>9 {x|x<-3 or x>3} zeros: ln(x^2-9)=0=ln(1) x^2-9=1 x^2=10 x=+sqrt10 or -sqrt10 inverse function: y=ln(x^2-9) e^y=x^2-9 x^2=e^y+9 x=+-sqrt(e^y+9) so we get a group of functions because f(x) is not injective one obvious possible function is f^-1(x)=sqrt(e^y+9) Rate Comment

OpenStudy (anonymous):

what is the x?

OpenStudy (anonymous):

meant f sorry